Big Mike’s, a large hardware store, has gathered data on its overhead activities and associated costs for the past 10 months. Nizam Sanjay, a member of the controller’s department, believes that overhead activities and costs should be classified into groups that have the same driver. He has decided that unloading incoming goods, counting goods, and inspecting goods can be grouped together as a more general receiving activity, since these three activities are all driven by the number of receiving orders. The 10 months of data shown below have been gathered for the receiving activity. Month Receiving Orders Receiving Cost 1 1,000 $12,170 2 1,340 12,940 3 1,150 13,750 4 900 9,930 5 1,350 15,070 6 1,400 14,145 7 1,600 16,640 8 1,490 14,800 9 1,800 17,940 10 1,700 15,000 Required: 1. Choose the correct scatter graph that represents the data provided above. a. b. c. d. The correct answer is c 2. Select two points (1 and 8) that make the best fit, and compute a cost formula for receiving costs. Round intermediate calculations to the nearest cent. Round the intercept to the nearest dollar and round the slope to the nearest cent. Y = $ + $X 3. Using the high-low method, prepare a cost formula for the receiving activity. Round the intercept to the nearest dollar and round the slope to the nearest cent. Y = $ + $X 4. Using the method of least squares, prepare a cost formula for the receiving activity. Round the intercept to the nearest dollar and round the slope to the nearest cent. Y = $ + $X What is the coefficient of determination? Round your answer to two decimal places. Adjusted R2 =
Answer #1

Part 2

Month

orders

cost

1

1000

12170

8

1490

14800

TC = Vx + F

14800=1490V + F

12170 = 1000V + F

2630 = 490V

V = 5.37

14800=(1490*5.37) + F

F = 14800-8001 = $6799

Therefore,

Y = $6799 + $5.37X

Part 3

Activity

orders

cost

Highest

1800

17940

Lowest

900

9930

TC = Vx + F

17940=1800V + F

9930 = 900V + F

8010 = 900V

V = 8.90

17940=(1800*8.90) + F

F = 17940-16020 = $1920

Therefore,

Y = $1920+ $8.90X

Part 4

Y = $4638+ $6.99X

Adjusted R2 = 0.80 or (79.84% in case of percentage form)

SUMMARY OUTPUT

Regression Statistics

Multiple R

0.905998

R Square

0.820832

Adjusted R Square

0.798436

Standard Error

1013.121

Observations

10

ANOVA

df

SS

MS

F

Significance F

Regression

1

37618885

37618885

36.65077

0.000305

Residual

8

8211318

1026415

Total

9

45830203

Coefficients

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Lower 95.0%

Upper 95.0%

Intercept

4637.984

1617.856

2.866748

0.020933

907.202

8368.766

907.202

8368.766

X Variable 1

6.992364

1.155001

6.053988

0.000305

4.328926

9.655802

4.328926

9.655802